Hydra Report - August 2002: Eng. Run

Updated on May 26, 2021, 11:18 am

More detector, dewar and Arcon work - Ramon Galvez , August 16th, 2002

Done:

  1. Tested the new macro thresholds set for cryo alarms. Working ok.
     
  2. With Rolo C. we made changes in the Hydra option of ‘SetFiles’. The waveforms and sequencer code are automatically installed now when Hydra + Arcon3.7 are selected.
     
  3. Tests were done also when switching back and forth for other ccds.
     
  4. With Ricardo V., we smoothly went through the LN2 solidification using the new device:

Normal operation vac sensors readouts

CCD dwr vac before pumping : 300 Torr
CCD dwr vac after pumping : 0 mT
   
LN2 flask vac after refill : 400 Torr
LN2 flask vac for solid state : 20 Torr ( after 1hr pumping )

 

To be done after this run (minor pending details) :

  1. Improve the ‘solid N2’ valve setpoint lock device.
     
  2. Install the new adjustable security ‘pop-up’ pressure valve.